OpenStudy (anonymous):

Simplify the expression: (3x^2y-5xy-12xy^2) (5xy^2+4xy)

OpenStudy (anonymous):

60y2 + 60y2 = 0 12x + 15xy + 0 = 20 + 73y + 60y2 12x + 15xy = 20 + 73y + 60y2